OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8

OpenStudy (anonymous):

So y is kx^2...

OpenStudy (anonymous):

since y=kx^2 and we know 54=k(3^2) therefore k=54/9=6

OpenStudy (anonymous):

384